You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121204 (25 Sep 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 3-4

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: M. Edwars
Instrument: LUCI1


At the beginning of the night there are thin cirrus close to the Orizon, wind 15m/s and a warning from temperature and dewpoint
The seeing is 1.1 at the beginning of Cresci_28A
We have pattern due to humidity in LUCI frames
Better seeing during the night ~0.8", but sometimes passing cirrus
Problems with LUCI imaging lamp flat: stange patter similar to dark frames
At the end of night we obtained LUCI lamp flats for Cresci + flats and arcs fot Hunt . Dark frames for Hunt can be found in UT20121205.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_28A DSF2237a H, 210_zJHK 1h Tell+flat. COMPLETED
Cresci_28A DSF2237a K, 210_zJHK 1h Tell+flat. Perhaps we have to repeat some few frames in K band due to passing cirrus.
Pentericci_02 Cluster1UDS HKspec, 200_H+K 20m ABORTED
Hunt_08 GRB020127 zJspec, 200_H+K 1h Tell+flat+arc. COMPLETED
Hunt_08 GRB020127 HKspec, 200_H+K 2h Tell+flat+arc. COMPLETED
Turatto_07 SN2012 J 200s imaging, no photometric standard required - COMPLETED
Turatto_07 SN2012 H 200s imaging, no photometric standard required - COMPLETED
Turatto_07 SN2012 Ks 200s imaging, no photometric standard required - COMPLETED

23.30 DoDarks.txt luci.20121204.0003-0049

00.13 We interrupt the darks to open the dome

00.55 Problems with the guider, the TO is trying to fix it.

01.25 We start with the Cresci_28A program.
  • Slew to the telluric Hipp104320
  • tell_acq_campoA.txt luci.20121204.0049-0055

01.38 tell_int_campoA_H.txt luci.20121204.0056-0063

01.50 tell_int_campoA_Koff.txt luci.20121204.0064-0067

01.53 we have to interrupt the science because Jhon Hill require to send one of the standard script that have to be ran at the beginning of the night

01.55 We slew to the target DSF2237a
  • acq_campoA.txt luci.20121204.0068-0076
  • we have some condensation in the II and III quadrants of LUCI FOV

02:22 cresci_campoA_H.txt luci.20121204.0077-00-88

03:24 cresci_campoA_K2.txt luci.20121204.0089-0100

04:32 Try to acquire tell_int_campoA_K.txt but target is too low

04:33 Check collimation (due to low elevation)

04:40 Slew to Pentericci1 telluric

04.45 pente1_telluric.acq luci.20121204.0101-0103

04,52 pente1_telluric.spec luci.20121204.0104-0112

05.03 aborted, the telluric is only inside the first slit while it should go around all the slits of the mask, the offsets of the script are wrong, we slew to science

05.05 pent1.acq luci.20121204.0113-118
  • No stars appears in lists, no reference boxes are available in the MOS masks
  • Try to perform an offset by hand in order to detect some object in the slit (x,y = 0.25,0)
  • Try again (x,y = 0.25,0)

05.35 pent1_HK_1.spec luci.20121204.0119-0122

05.59 We need to close the dome due to high humidity

06.40 Opening dome

6.50 Problem with the secondary mirror, the TO is trying to fix it.

07.04 GRB02127acquisition_tell.script luci.20121204.0123-0125

07.17 GRB020127science_zJspec_tell.script luci.20121204.0126-0134
  • thin cirrus

07.28 GRB020127acquisition.script luci.20121204.0135-0138

07.51 GRB020127science_zJspec.script luci.20121204.0139-144

08:55 GRB02127acquisition_tell.script luci.20121204.0145-147

09:02 GRB020127science_zJspec_tell.script luci.20121204.0148-0156

09:16 GRB020127science_HKspec_tell.script ==luci.20121204.0157-0165

09:27 GRB020127acquisition.script ==luci.20121204.0166-0169

09.40 GRB020127science_HKspec.script luci.20121204.0170-193
  • good seeing during the integration 0.8"
  • cirrus transit

11:49 GRB02127acquisition_tell.script luci.20121204.0194-0196

11:55 GRB020127science_HKspec_tell.script luci.20121204.0197-0205

12:07 Changing camera to acquire Turatto SN2012ab

12:15 SN2012ab_J.img luci.20121204.0206-0209
  • first pointing failed. We chance the guiding star

12:40 SN2012ab_H.img luci.20121204.0210-0213

12:46 SN2012ab_Ks.img luci.20121204.021-0217

13.00 Dome closed

13.03 lamp flat started, J band, comand flatJ.txt =luci.20121204.0220 =
  • not good because the calibration unit was not in

13.05 lamp flat started, J band, comand flatJ.txt luci.20121204.0221-0228==
  • lamp flats show a strange pattern, similar to darks. At the moment, no explanation for this. We will interact with D. Thomson about this issue.

13.10 We start the flat and arc integrations for Hunt_08 and repeat Cresci_28 to take under control the humidity features.
  • flat_campoA_K.txt luci.20121204.0229-0238
  • flat_campoA_H.txt luci.20121204.0239-0248
  • grbhosts_HKflat.script luci.20121204.0249-0268
  • grbhosts_zJflat.script luci.20121204.0269-0???
  • grbhosts_zJarc.script luci.20121204.03???-0300
  • grbhosts_HKarc.script luci.20121204.0301-??

-- VincenzoTesta - 04 Dec 2012
Topic revision: r13 - 25 Sep 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback